2x-3y=-6
3x+4y=12














Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Did you bother to read the instructions on the page where you go to post your questions?

Here they are again if you either missed them, didn't understand them, or simply ignored them.

Read the Instructions

John

My calculator said it, I believe it, that settles it


Answer by stanbon(75887)

2x-3y=-6
3x+4y=12
----------------
Multiply thru 1st by 3
Multiply thru 2nd by 2
----------------
6x - 9y = -18
6x + 8y = 24
----------------
Subtract 1st from 2nd and solve for "y":
17y = 42
y = 42/17
-------------
Substitute for "y" and solve for "x":
3x + 4(42/17) = 12
3x + (168/17) = 204/17
3x = 36/17
x = 12/17
===================================
Cheers,
Stan H.